# Generalized saddle-node ghosts and their composite structures in dynamical systems

### 09 Apr 2026 v2

@@ 2.4.2 Timescale induced transitions in ghosts of coupled climate tipping elements
where $x_{i}$ denotes the current value of the $i$th tipping element, $\tau_{i}$ is its tipping timescale, $\Delta GMT$ is the increase in the global mean surface temperature above pre-industrial levels, $T_{\text{limit},i}$ is the critical temperature above which the element tips (passes a SN-bifurcation), $d$ is an overall interaction strength parameter and $s_{i,j}$ represents the link strength between two interacting tipping elements. Since Equation 7 explicitly features multiple timescales, it is also an interesting system to explore how slow-fast dynamics and ghost dynamics interact. While a detailed investigation of the ghost structures in this system is the subject of another study [Nandan et al., in preparation], we focus here on two generic tipping elements which mutually reinforce each other. We assume the two tipping elements have similar but not identical critical temperatures and that their tipping timescales are different by a factor of 10. Generating several trajectories that sample the phase space using PyGhostID’s ghostID_\_phaseSpaceSample function, we identify three ghosts of dimension 1 for $\Delta GMT$ being set just above the highest value of $T_{limit,i}$ (Figure 5a). Following a trajectory starting at $x_{1}(0)=-1.5,x_{2}(0)=-1.2$, we see the trajectory visiting both ghosts $G3$ and $G1$ before arriving at a stable fixed point (Figure 5a, white line), suggesting the system exhibits a ghost channel [38] between G3 and G1. To confirm this, we use the ghost_\_connections function on the ensemble of trajectories to generate an adjacency matrix of ghost-ghost sequences. In addition to plotting the corresponding graph, we can encode information about each ghost captured by GhostID visually via the node color, e.g., to represent the trapping duration at each ghost (Figure 5b). To confirm that ghostID_\_phaseSpaceSample identified all ghosts, we further followed each ghost versus varying $\Delta GMT$ values using track_\_ghost_\_branch, showing that each ghost branch terminates at one of three SN-bifurcation points in which a stable fixed point and a saddle merge (Figure 5c).
@@
Next, we explored the influence of tipping timescales in this context. We thus characterize the system’s ghost dynamics as we systematically increase $\tau_{1}$ until $\tau_{1}>\tau_{2}$. As track_\_ghost_\_branch only tracks a single ghost, we used ghostID_\_phaseSpaceSample in combination with ghost_\_connections to identify the dimensions and the connections between the ghosts. While two of the ghosts exhibited no changes, we found a transition in the dimension of G3 from 1 to 2 at $\tau_{1}\approx 1.5\ \tau_{2}$ (Figure 5d, upper panel). Interestingly, considering the number of ghost connections as quantified by the adjacency matrix reveals another transition: as $\tau_{1}$ increases, the number of ghost connections drops to 0 at $\tau_{1}=3500$ before rising to 1 again at $\tau_{1}\geq 18500$ (Figure 5d, lower panel). Visualizing the corresponding graphs reveals that these transitions are associated with a re-wiring of the ghost connections. While for low values of $\tau_{1}$, we have a ghost channel from G3 to G1, a new ghost channel from to G3 to G2 is formed for high $\tau_{1}$ values. While these two types of transitions are clearly different from each other as indicated by different critical thresholds, they are also different from the ghost SN-bifurcation described above, in which an attracting and a non-attracting ghost merge. Zooming in into the phase space region around G3, we can see how increasing $\tau_{1}$ changes both the distribution of eigenvalues across space and the direction of the flow around G3 (Figure 5e). At low $\tau_{1}$-values, we see only eigenvalue $\lambda_{1}$ around G3 being distributed from negative to positive along the flow, while $\lambda_{2}$ stays purely negative around G3. As $\tau_{1}$ increases, the flow compressed in G3 is more spread in the outbound region of G3 where now $\lambda_{2}$, too, becomes positive, thereby allowing trajectories to sample $\lambda_{1}$ from negative to positive. Hence, the change in the dimension of G3 is as result of both the changed eigenvalue distribution across space and of the altered flow, which together determine what eigenvalues a trajectory will see as it passes close to or through G3. The transition between ghost connections, on the other hand, is a result only of the altered flow that determines if a another ghost can be reached starting from G3 or not.
@@ 2.4.3 Ghost networks in gene regulatory networks at criticality
where $x_{i}$ refers to the concentration of gene product $i$, $\alpha$ is the auto-activation strength, $\beta$ the maximum transcription rate of activators, $a_{ij}$ are the elements of the adjacency matrix AA that describes the GRN’s interactions, and $K,K_{a},K_{i}$ are the half-activation constants of (auto-)activation and repression, respectively. For $K=0.5$ and in the absence of any activators and inhibitors, each gene exhibits a SN-bifurcation at $\alpha=1$ (Figure 6a). We assume here $\alpha=0.998$, i.e., each gene is organized close to the SN-bifurcation and may thus exhibit a simple ghost, not unlike the situation described in [20]. Directed Erdős–Rényi networks for which each activatory link was replaced with an inhibitory link with a probability of $0.5$ were used for to generate different GRN topologies.
@@
Using this simple network model, we observed larger ghost structures including ghost channels, ghost cycles and various combinations thereof, with many ghosts having a dimension larger than one. As an example we show a network with $N=12$ nodes (Figure 6b). Using a large phase space sample of this network generated by ghostID_\_phaseSpaceSample, we identified the ghost structures embedded in the GRN using the ghost_\_connections function and colored each node in the corresponding graph according to the ghost dimensions. As shown in Figure 6c, we find a strikingly complex ghost network embedded in the phase space of the GRN, featuring many interesting motifs, including cyclical structures, several ghosts converging into one, and ghost ejecting their flow to two or more other ghosts (e.g. G6). While a systematic characterization of such ghost networks in phase space and their relation to the network structure of the dynamical system is beyond the scope of the current study, a few interesting observations can be drawn from this example already. First, the ghost network has a different topology and is notably larger than the GRN, featuring both more nodes and edges. The topology of the ghost network further suggests multi-stability between a three-ghost cycle, at which the majority of nodes eventually converge by following their links, and two ghost channels from G5 and G9 to G10 after which trajectories converge at an attractor. Time courses of trajectories starting at three random initial conditions confirm this, showing convergence to the three ghost cycle with its periodic switching between distinct long transients of gene expression after different initial transients for two initial conditions and a two-ghost channel that ends at a stable steady state for the third initial conditions (Figure 6d). Another interesting observation is that the dimensions of the ghosts in the ghost network are relatively small compared to the dimension of the system ($n=12$), with the majority of ghosts having a dimension of 2-4. The dimension of the ghosts furthermore do not necessarily correspond to the number of connections a ghost can form. For example, while G9 has a ghost dimension of 4, it only connects to two other ghosts. Sampling initial conditions around this ghost and projecting the high-dimensional trajectories into a three-dimensional principle component space shows that trajectories are being spread into four different directions consistent with the ghost dimension measured by GhostID before two pairs of the four trajectory bundles are each being funneled towards G4 and G10, respectively (Figure 6e).
@@ 3 Discussion
We have presented here a generalized definition of ghosts in dynamical systems that accounts for the diverse topological structures of the saddle-node fixed points they are the remnant of and developed a series of algorithms to identify and characterize ghosts and their composite structures such as ghost channels and ghost cycles [38]. Naturally, there are several open questions and limitations that come with our approach. First, our current framework was developed to capture ghosts of fixed points. While further work is necessary to see whether it applies to ghosts of continuous attractors [23, 56], the current algorithms cannot capture ghosts of limit cycles as the slow dynamics only apply along the radial direction of their periodic orbit and it is unclear if any meaningful eigenvalue characteristics can be extracted from ghosts of limit cycles. The same argument applies to ghosts from chaotic attractors [44, 9]. At least the first problem has been solved by the elegant study from Zheng et al. [77], who used variational methods to create a ‘best-fit’ topological structure in the vector field. An alternative approach may be to replace the scalar function $Q$ by recurrences [15] for finding non-invariant structures in phase space that a trajectory spends a long time in. Secondly, the framework was designed for autonomous deterministic systems. While we suspect that slow parameter drifts will not impede detection of ghosts by our algorithm, faster or more complex parameter changes in non-autonomous systems will likely interfere with the algorithm. Similar considerations apply to stochastic systems. Noise has a significant influence on a trajectories’ dwell-time at a ghost and can push it back and forth along the ghost region [57, 38] which may result in non-monotonic eigenvalue samples that impair ghost identification. Lastly, we assumed throughout that $f$ is analytic. While this assumption holds for a vast number of models from applications, additional assumptions are likely required for our framework to apply smooth systems (cf. Supplementary section 1.7), although these assumptions are unlikely to affect the here introduced algorithms in most applications.
@@
Despite these limitations we showed that our definitions and algorithms offer a powerful new approach to study ghost dynamics in arbitrary dynamical systems. Using several relatively simple example systems, this approach allowed us to discover several interesting phenomena related to ghosts. While their comprehensive characterization is beyond the scope of the current study, we will briefly discuss their implications.
@@
In our first example, two coupled and identical theta neurons [5], we showed the existence of a two-dimensional ghost following the merging of a sink, a repeller and two saddles (a phase portrait sometimes called basic tartan [6]) at a type $2,0$ saddle-node of a degenerate SNIC-bifurcation. Since the theta neuron model is the normal form of a SNIC bifurcation, there is strong reason to believe that this result will apply to many coupled systems whose units individually undergo SNIC bifurcations. Characterizing the ghost dynamics in the case of non-identical neurons with different excitability thresholds, we found that one- and two-dimensional ghosts are detectable across a large area of the parameter plane, although pronounced long transients are only found close to bifurcation. This two-parameter analysis further revealed the existence of what we termed a ghost SN-bifurcation, a collision of two one-dimensional ghosts - one of which is attracting, one non-attracting in the sense of definition 2.2 - leading to the formation of a two-dimensional ghost. To our knowledge, this is the first description of a bifurcation that occurs exclusively between non-invariant objects and produces another non-invariant object. From a biological perspective, the example raises interesting new questions about the role of ghost dynamics in the context of information processing and computations in neuronal networks. For example, could higher dimensional ghosts provide a reliable mechanism for decoding multi-modal time-varying inputs as compared to single time-varying inputs considered previously [13]? Assuming larger ghost networks as found in our GRN example can also emerge from neuronal dynamics, another intriguing perspective is to consider ghosts and their composite structure as a potential transient-based mechanism for structuring phase space trajectories along neuronal manifolds [50]. As we find the properties of ghosts to be tunable by control parameters, it is tempting to speculate that neuronal networks might be able to adapt the properties of individual ghosts and larger ghost structures to their computational needs. Studying how ghost dynamics structure and support biological computations is thus an important area for further research.
@@
In the second example, we examined coupled climate tipping elements using the model from Wunderling et al. [73]. We focused here on two the case of two generic, mutually reinforcing tipping elements with similar temperature thresholds but different tipping timescales in a scenario when the global mean temperature just passed these thresholds. Our analysis revealed the existence of three ghost states in this model, two of which form a ghost channel, detectable for the whole range of global warming considered here (i.e., up to $\Delta GMT=3$°C). We further found a striking influence of the tipping timescales on the system’s ghost dynamics, which included the change of one of the ghosts dimension by altering the local flow and eigenvalue distribution as well as a restructuring of the ghost connections. Since neither of these transition involved a collision between ghosts, the observed transitions are distinct from the ghost SN-bifurcation observed in the coupled theta neuron model. However, further work will be necessary to better understand the nature of the timescale-dependent transitions. Since long transients due to ghosts can keep a system in a desirable area of phase space even when the equilibrium is lost, a more detailed investigation of the ghost dynamics in a higher dimensional version of this model for realistic parameters, real-world network topologies and policy scenarios is currently underway [Nandan et al., in preparation].
@@
In the final example, we applied our algorithms to GRNs under the assumption that each gene is poised at criticality, i.e. close to a SN-bifurcation. Our analysis showed that embedded in the dynamics of such GRNs can be a complex ghost network with a topology distinct from the GRN and many higher-dimensional ghosts. These ghost networks can further be multi-stable in the sense of coexisting ghost structures, such that some initial conditions lead to visitation, e.g., of a ghost channel, whereas others end up on a ghost cycle. While the assumption of each gene being poised at criticality is probably too strong to be reflective of real GRNs, there is a growing body of empirical evidence for ghost dynamics in GRNs [53, 45] and intracellular signaling networks[62]. While the role of ghost dynamics in cell biology is barely explored, its unique properties have been used to explain processing of time-varying stimuli[63, 37] and put forward as robust mechanisms for cell population control[61], regulation of developmental transitions[53], and cell-type specification and differentiation[45, 20]. Moreover, our example demonstrates that complex ghost dynamics such as described in [20] can easily emerge in random networks and do not require artificial network structures. Exploring complex and higher-dimensional ghost dynamics in GRNs and other types of intracellular regulatory networks thus has the potential provide new perspectives on various phenomena in cell biology.
@@
Based on these results we believe the framework and algorithms presented here will open many interesting avenues for further research ranging from physical sciences to living systems.

### 16 Apr 2026 v3

@@ 2.4.2 Timescale induced transitions in ghosts of coupled climate tipping elements
where $x_{i}$ denotes the current value of the $i$th tipping element, $\tau_{i}$ is its tipping timescale, $\Delta GMT$ is the increase in the global mean surface temperature above pre-industrial levels, $T_{\text{limit},i}$ is the critical temperature above which the element tips (passes a SN-bifurcation), $d$ is an overall interaction strength parameter and $s_{i,j}$ represents the link strength between two interacting tipping elements. Since Equation 7 explicitly features multiple timescales, it is also an interesting system to explore how slow-fast dynamics and ghost dynamics interact. While a detailed investigation of the ghost structures in this system is the subject of another study [Nandan et al., in preparation], we focus here on two generic tipping elements which mutually reinforce each other. We assume the two tipping elements have similar but not identical critical temperatures and that their tipping timescales are different by a factor of 10. Generating several trajectories that sample the phase space using PyGhostID’s ghostID_\_phaseSpaceSample function, we identify three ghosts of dimension 1 for $\Delta GMT$ being set just above the highest value of $T_{\text{limit},i}$ (Figure 5a). Following a trajectory starting at $x_{1}(0)=-1.5,x_{2}(0)=-1.2$, we see the trajectory visiting both ghosts G3 and G1 before arriving at a stable fixed point (Figure 5a, white line), suggesting the system exhibits a ghost channel [38] between G3 and G1. To confirm this, we use the ghost_\_connections function on the ensemble of trajectories to generate an adjacency matrix of ghost-ghost sequences. In addition to plotting the corresponding graph, we can encode information about each ghost captured by GhostID visually via the node color, e.g., to represent the trapping duration at each ghost (Figure 5b). To confirm that ghostID_\_phaseSpaceSample identified all ghosts, we further followed each ghost while varying $\Delta GMT$ values using track_\_ghost_\_branch, showing that each ghost branch terminates at one of three SN-bifurcation points in which a stable fixed point and a saddle merge (Figure 5c).
@@
Next, we explored the influence of tipping timescales in this context. We thus characterize the system’s ghost dynamics as we systematically increase $\tau_{1}$ until $\tau_{1}>\tau_{2}$. As track_\_ghost_\_branch only tracks a single ghost, we used ghostID_\_phaseSpaceSample in combination with ghost_\_connections to identify the dimensions and the connections between the ghosts. While two of the ghosts exhibited no changes, we found a transition in the dimension of G3 from 1 to 2 at $\tau_{1}\approx 1.5\ \tau_{2}$ (Figure 5d, upper panel). Interestingly, considering the number of ghost connections as quantified by the adjacency matrix reveals another transition: as $\tau_{1}$ increases, the number of ghost connections drops to 0 at $\tau_{1}=3500$ before rising to 1 again at $\tau_{1}\geq 18500$ (Figure 5d, lower panel). Visualizing the corresponding graphs reveals that these transitions are associated with a re-wiring of the ghost connections. While for low values of $\tau_{1}$, there is a ghost channel from G3 to G1, a new ghost channel from to G3 to G2 is formed for high $\tau_{1}$ values. While these two types of transitions are clearly different from each other as indicated by different critical thresholds, they are also different from the ghost SN-bifurcation described above, in which an attracting and a non-attracting ghost merge. Zooming in into the phase space region around G3, we can see how increasing $\tau_{1}$ changes both the distribution of eigenvalues across space and the direction of the flow around G3 (Figure 5e). At low $\tau_{1}$-values, we see only eigenvalue $\lambda_{1}$ around G3 being distributed from negative to positive along the flow, while $\lambda_{2}$ stays purely negative around G3. As $\tau_{1}$ increases, the flow compressed in G3 is more spread in the outbound region of G3 where now $\lambda_{2}$, too, becomes positive, thereby allowing trajectories to sample $\lambda_{1}$ from negative to positive. Hence, the change in the dimension of G3 is as result of both the changed eigenvalue distribution across space and of the altered flow, which together determine what eigenvalues a trajectory will see as it passes close to or through G3. The transition between ghost connections, on the other hand, is a result only of the altered flow that determines if a another ghost can be reached starting from G3 or not.
@@ 2.4.3 Ghost networks in gene regulatory networks at criticality
where $x_{i}$ refers to the concentration of gene product $i$, $\alpha$ is the auto-activation strength, $\beta$ the maximum transcription rate of activators, $a_{ij}$ are the elements of the adjacency matrix $A$ that describes the GRN’s interactions, and $K,K_{a},K_{i}$ are the half-activation constants of (auto-)activation and repression, respectively. For $K=0.5$ and in the absence of any activators and inhibitors, each gene exhibits a SN-bifurcation at $\alpha=1$ (Figure 6a). We assume here $\alpha=0.998$, i.e., each gene is organized close to the SN-bifurcation and may thus exhibit a simple ghost, not unlike the situation described in [20]. Directed Erdős–Rényi networks for which each activatory link was replaced with an inhibitory link with a probability of $0.5$ were used to generate different GRN topologies.
@@
Using this simple network model, we observed larger ghost structures including ghost channels, ghost cycles and various combinations thereof, with many ghosts having a dimension larger than one. As an example we show a network with $N=12$ nodes (Figure 6b). Using a large phase space sample of this network generated by ghostID_\_phaseSpaceSample, we identified the ghost structures embedded in the GRN using the ghost_\_connections function and colored each node in the corresponding graph according to the ghost dimensions. As shown in Figure 6c, we find a strikingly complex ghost network embedded in the phase space of the GRN, featuring many interesting motifs, including cyclical structures, several ghosts converging into one, and ghosts ejecting their flow to two or more other ghosts (e.g. G6). While a systematic characterization of such ghost networks in phase space and their relation to the network structure of the dynamical system is beyond the scope of the current study, a few interesting observations can be drawn from this example already. First, the ghost network has a different topology and is notably larger than the GRN, featuring both more nodes and edges. The topology of the ghost network further suggests multi-stability between a three-ghost cycle, at which the majority of nodes eventually converge by following their links, and two ghost channels from G5 and G9 to G10 after which trajectories converge at an attractor. Timecourses of trajectories starting at three random initial conditions confirm this, showing convergence to the three ghost cycle with its periodic switching between distinct long transients of gene expression after different initial transients for two initial conditions and a two-ghost channel that ends at a stable steady state for the third initial condition (Figure 6d). Another interesting observation is that the dimensions of the ghosts in the ghost network are relatively small compared to the dimension of the system ($n=12$), with the majority of ghosts having a dimension of 2-4. The dimension of the ghosts furthermore do not necessarily correspond to the number of connections a ghost forms. For example, while G9 has a ghost dimension of four, it only connects to two other ghosts. Sampling initial conditions around this ghost and projecting the high-dimensional trajectories into a three-dimensional principle component space shows that trajectories are being spread into four different directions, consistent with the ghost dimension measured by GhostID, before two pairs of the four trajectory bundles are each being funneled towards G4 and G10, respectively (Figure 6e).
@@ 3 Discussion
We have presented here a generalized definition of ghosts in dynamical systems that accounts for the diverse topological structures of the saddle-node fixed points they are the remnant of and developed a series of algorithms to identify and characterize ghosts and their composite structures such as ghost channels and ghost cycles [38]. Naturally, there are several open questions and limitations that come with our approach. First, our current framework was developed to capture ghosts of fixed points. While further work is necessary to see whether it applies to ghosts of continuous attractors [23, 56], the current algorithms cannot capture ghosts of limit cycles as the slow dynamics only apply along the radial direction of their periodic orbit and it is unclear if any meaningful eigenvalue characteristics can be extracted from ghosts of limit cycles. The same argument applies to ghosts from chaotic attractors [44, 9]. At least the first problem has been solved by the elegant study from Zheng et al. [77], who used variational methods to create a ‘best-fit’ topological structure in the vector field. An alternative approach may be to replace the scalar function $Q$ by recurrences [15] for finding non-invariant structures in phase space that a trajectory spends a long time in. Secondly, the framework was designed for autonomous deterministic systems. While we suspect that slow parameter drifts will not impede detection of ghosts by our algorithm, faster or more complex parameter changes in non-autonomous systems will likely interfere with the algorithm. Similar considerations apply to stochastic systems. Noise has a significant influence on a trajectories’ dwell-time at a ghost and can push it back and forth along the ghost region [57, 38] which may result in non-monotonic eigenvalue samples that impair ghost identification. Lastly, we assumed throughout that $f$ is analytic. While this assumption holds for a vast number of models from applications, additional assumptions are likely required for our framework to apply to smooth systems (cf. Supplementary section 1.7), although these assumptions are unlikely to affect the here introduced algorithms in most cases.
@@
Despite these limitations, we showed that our definitions and algorithms offer a powerful new approach to study ghost dynamics in arbitrary dynamical systems. Using several relatively simple example systems, this approach allowed us to discover several interesting phenomena related to ghosts. While their comprehensive characterization is beyond the scope of the current study, we will briefly discuss their implications.
@@
In our first example, two coupled and identical theta neurons [5], we showed the existence of a two-dimensional ghost following the merging of a sink, a repeller and two saddles (a phase portrait sometimes called basic tartan [6]) at a type $2,0$ saddle-node of a degenerate SNIC-bifurcation. Since the theta neuron model is the normal form of a SNIC bifurcation, there is strong reason to believe that this result will apply to many coupled systems whose units individually undergo SNIC bifurcations. Characterizing the ghost dynamics in the case of non-identical neurons with different excitability thresholds, we found that one- and two-dimensional ghosts are detectable across a large area of the parameter plane, although pronounced long transients are only found close to bifurcation. This two-parameter analysis further revealed the existence of what we termed a ghost SN-bifurcation, a collision of two one-dimensional ghosts - one of which is attracting, one non-attracting in the sense of definition 2.2 - leading to the formation of a two-dimensional ghost. To our knowledge, this is the first description of a bifurcation that occurs exclusively between non-invariant objects and produces another non-invariant object. From a biological perspective, the example raises interesting new questions about the role of ghost dynamics in the context of information processing and computations in neuronal networks. For example, could higher dimensional ghosts provide a reliable mechanism for decoding multi-modal time-varying inputs as opposed to single time-varying inputs considered previously [13]? Assuming larger ghost networks as found in our GRN example can also emerge from neuronal dynamics, another intriguing perspective is to consider ghosts and their composite structure as a potential transient-based mechanism for structuring phase space trajectories along neuronal manifolds [50]. As we find the properties of ghosts to be tunable by control parameters, it is tempting to speculate that neuronal networks might be able to adapt the properties of individual ghosts and larger ghost structures to their computational needs. Studying how ghost dynamics structure and support biological computations is thus an important area for further research.
@@
In the second example, we examined coupled climate tipping elements using the model from Wunderling et al. [73]. We focused here on two the case of two generic, mutually reinforcing tipping elements with similar temperature thresholds but different tipping timescales in a scenario when the global mean temperature just passed these thresholds. Our analysis revealed the existence of three ghost states in this model, two of which form a ghost channel, detectable for the whole range of global warming considered here (i.e., up to $\Delta GMT=3$°C). We further found a striking influence of the tipping timescales on the system’s ghost dynamics, which included the change of one of the ghosts dimension by altering the local flow and eigenvalue distribution as well as a restructuring of the ghost connections. Since neither of these transition involved a collision between ghosts, the observed transitions are distinct from the ghost SN-bifurcation observed in the coupled theta neuron model. However, further work is required to better understand the nature of the timescale-dependent transitions. Since long transients due to ghosts can keep a system in a desirable area of phase space even when the equilibrium is lost, a more detailed investigation of the ghost dynamics in a higher dimensional version of this model for realistic parameters, real-world network topologies and policy scenarios is currently underway [Nandan et al., in preparation].
@@
In the final example, we applied our algorithms to GRNs under the assumption that each gene is poised at criticality, i.e., close to a SN-bifurcation. Our analysis showed that embedded in the dynamics of such GRNs can be a complex ghost network in phase space with many higher-dimensional ghosts and a topology distinct from the GRN. These ghost networks can further be multi-stable in the sense of coexisting ghost structures, such that some initial conditions lead to visitation, e.g., of a ghost channel, whereas others end up on a ghost cycle. While the assumption of each gene being poised at criticality is perhaps too strong to be representative of real GRNs, there is a growing body of empirical evidence for ghost dynamics in GRNs [53, 45] and intracellular signaling networks [62]. While the role of ghost dynamics in cell biology is barely explored, its unique properties have been used to explain processing of time-varying stimuli[63, 37] and put forward as robust mechanisms for cell population control[61], regulation of developmental transitions [53], and cell-type specification and differentiation [45, 20]. Moreover, our example demonstrates that complex ghost dynamics such as described in [20] can easily emerge in random networks and do not require artificial network structures. Exploring complex and higher-dimensional ghost dynamics in GRNs and other types of intracellular regulatory networks thus has the potential to provide new perspectives on various phenomena in cell biology.
@@
Based on these results, we believe the framework and algorithms presented here will open many interesting avenues for further research ranging from physical sciences to living systems.